OpenStudy (anonymous):

Write an equation in slope-intercept form of the line through points S(–1, –4) and T(3, 4)

OpenStudy (anonymous):

Ok this is easy to do. You need two things to make an equation: a point and the slope.

OpenStudy (anonymous):

We have two points, so we can only either one. Now for the slope, we need to use the formula for finding slope. Do you remember the formula 

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Ok so we plug in the points. Let S be (x2,y2) and let T be (x1, y1)

OpenStudy (anonymous):

we can also let S be (x1, y1) and T be (x2,y2). It doesn't matter

OpenStudy (anonymous):

I'll let S be (x1, y1) and T be (x2, y2). Ok so all we have to do is plug in these numbers into the formula and see what the slope is.

OpenStudy (anonymous):

Plug them in, then solve the slope and tell me what you get.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Plug in -1 for x1, -4 for y1, 3 for x2, and 4 for y2.

OpenStudy (anonymous):

4+4 ---- 3+1 right?

OpenStudy (anonymous):

Yup, and that simplifies to 2.

OpenStudy (anonymous):

We now have our slope, and at least one point on the line. We have all we need to find the equation.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Do you remember the point-slope formula?

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

Yup so just plug in 2 for m.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

For x1 and y1, you can use EITHER (3,4) or (-1,-4). Tip: Don't take the x-value of one and the y-value of the another. For example don't do (3,-4).

OpenStudy (anonymous):

So let's try it with (3,4). Our x1 is 3 in this case and our y1 is 4.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Try plugging these into y-y1= m(x-x1) and see what you get.

OpenStudy (anonymous):

(y-3)=2(x-4)

OpenStudy (anonymous):

Great, so all we have to do now is put the equation into slope-intercept form

OpenStudy (anonymous):

Do you know how to do that?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

no Thats what im struggling with the most

OpenStudy (anonymous):

Ok so first of all. Do you see the part 2(x-4)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

We use the distributive property. Do you know how to do it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2x-4

OpenStudy (anonymous):

Well that's almost correct

OpenStudy (anonymous):

is it y=2x-2 ?

OpenStudy (anonymous):

Yup the answer is y = 2x-2

OpenStudy (anonymous):

But when you do the distributive rule, you shouldn't get 2x-4

OpenStudy (anonymous):

You should learn this as this will come up in math for your entire high school

OpenStudy (anonymous):